OpenStudy (anonymous):

Will give medals! What is an equation in standard form of an ellipse centered at the origin with vertex (-5,0) and co-vertex (0,3)?

OpenStudy (anonymous):

x^2+y^2=25

OpenStudy (anonymous):

is this a general equation I have to fill in?? My answer options are: A. (X^2/25) + (Y^2/9)=1 B. (X^2/5) + (Y^2/3)=1 C. (X^2/9) + (Y^2/25)=1 D. (X^2/3) + (Y^2/5)=1

OpenStudy (anonymous):

i believe it would be C

OpenStudy (anonymous):

thank you!